Two events  and  associated with a sample space of an experiment are such that        ,  and  . Determine whether   and  are independent.

(b)      Two events  and  are such that, they are independent and  and

             and  . Find

i)           the value of

ii)           and  , where  are the complements of events  and  respectively.


1
Expert's answer
2021-07-20T17:25:31-0400

(a) Given "E_1" and "E_2" are the events that are associated with a sample space of experiment.


"P(E_1) = 0.27\\

P(E_2) = 0.40\\

P(E_1\cup E_2)=0.58"


Condition: For two events "E_1" and "E_2" are independent iff

"P(E_1\cap E_2)=P(E_1)\cdot P(E_2)" , otherwise they are not independent

So, by addition law of probability

"P(E_1\cap E_2)=P(E_1)+P(E_2)-P(E_1\cup E_2)\\"

"=0.27+0.40-0.58\\=0.09"


and "P(E_1)\cdot P(E_2)=0.27\times 0.40=0.108"


So, we can observe that "P(E_1\cap E_2)\neq P(E_1)\cdot P(E_2)"

Hence, we can say that "E_1" and "E_2" are not independent events.


(b) Given , two events A and B , such that they are independent

"P(A)=x\\P(B)=x+0.2\\P(A\cap B)= 0.15"


If A and B are independent events, then

"P(E_1\cap E_2)= P(E_1)\cdot P(E_2)\\0.15=x(x+0.2)\\20x^2+4x-3=0"

Upon solving the quadratic equation,

we get "\boxed{x=\dfrac{3}{10}=0.3}" and x = -0.5 ("x \neq -0.5" must have a positive value)


option (i) and (ii is not clearly shown in the problem. )